# Wearable Roller Rings to Enable Robot Dexterous In-Hand Manipulation through Active Surfaces

### v1

@@ IV-B Multi-Sphere Motion Model for Arbitrary Object Shapes
Extending from the rotation motion model on a unit sphere, we now model the rotations of arbitrary object shapes as actuated by active surface contacts. Let us still begin with the scenario with $2$ contacts $p_{a}$ and $p_{b}$ on an object. However, this time the contacts are not necessarily on a unit sphere $\mathbb{S}^{2}$. Their distances to the axis of rotation are represented by $r_{a},r_{b}\in\mathbb{R}$. Intuitively, these two contacts are actuating the object as if there were $2$ different spheres for them to make contacts, as illustrated in Fig. 5. As such, the torques generated by these two contacts, as shown in Eq. (9), are now additionally scaled by $r_{a},r_{b}$:
@@
As such, similarly to the derivation for the motion model on a unit sphere, the torques should all balance out when the object is rotating at an equilibrium state. Skipping the intermediate steps, we can obtain the rotation motion model for $N$ active surface contacts on arbitrary objects:
@@
Based on this multi-sphere motion model, Eq. (10) also provides a complete and non-holonomic rotational manipulation solution. However, a major difference now is that the individual angular velocities are additionally weighted by the distances between the contact and the rotation axis, providing another dimension for rotations to be regulated by contact variances.
@@ IV-C Motion Model for In-hand Translations
To achieve a complete in-hand manipulation solution, the final component required in our motion model is the ability to translate the object in-hand. Our model analysis here will focus on pure translations supplementary to the rotation models derived above. In practice, however, the object rotation and translation can happen simultaneously during manipulation in order to perform more efficient actions.
For pure in-hand object translation, the first requirement is that the object’s angular velocity should be zero. Based on our model above, we know this can be expressed as a motion constraint among all active surfaces: $\omega_{*}=\sum_{i=1}^{N}r_{i}k_{i}\omega_{i}=0$. When this constraint is satisfied, which can be achieved in an infinite number of ways when multiple contacts exist to cancel out each other’s torques, the velocity of the object is purely determined by the linear forces provided by the active surfaces. Let us denote by $v^{*}\in\mathbb{R}^{3}$ the linear velocity of the object, by $v_{i}\in\mathbb{R}^{3}$ and $F_{i}^{t}$ the linear velocity and the tangential force at the $i$-th active surface. The following equation further constrains the object’s translation:
@@
In Eq. (11), the tangential force at each contact is first aligned to the direction of the “slipping or scratching” motion and then projected to the perpendicular direction of the object’s motion (by dot product). The zero-sum of all these forces guarantees that the object is purely translating in the direction of $v^{*}$. In practice, although it is not difficult to obtain $\omega_{*}=0$ as discussed above, achieving arbitrary translation motions is challenging due to the additional constraint (Eq. (11)), especially when the number of contacts is small. Nonetheless, given our complete solution of in-hand rotation, we will be able to manipulate the object from any pose to any others by sequencing translation and rotations. As shown in [28], a manipulation graph can be constructed to generate such action sequences, although the aforementioned “detours” are not avoidable.
@@ V Experiments
Figure 6: Experimental Objects manipulated using the Roller Rings with the Yale Model O (Top Row Left to Right: A-F Sphere, Octahedra, Children’s Cheese Toy, Cardboard Tube; Bottom Row Left to Right: A-F Sphere, Cube, Children’s Grape Toy)
Figure 7: Motion Examples of the RR: (a)A-F Sphere reoriented by the RR through all faces when held in a power grasp by the Yale Model O (b) Cardboard tube translated along the Yale Model O when held in a power grasp
To evaluate the design and motion model of the RR, we constructed a prototype, as seen in Fig. 2(a) to test its in-hand manipulation capabilities using various objects (Fig. 6). The RRs were affixed to both a Yale Model O hand and a human hand for all experiments wherein all experiments were conducted using teleoperation. As shown in Sec. III, each RR variant has one actuated DoF from the Micro DC Motor that was controlled using velocity control. Each RR had its velocity set to a differing set of values during the experiment so that the combinations of varying velocities can generate differential motions for the grasped object, thus allowing for a complete manipulation solution (Sec. IV).
@@ V-A Rotational Manipulation
While conducting the experiments, we evaluated the range of rotation that was physically possible by the RRs through their combined set of actuated DoFs. In this, so long as the manipulated object’s geometry didn’t collide internally with the grasping system and a stable grasp was maintained, we aimed to confirm that the designed RRs could rotate a grasped object from any initial orientation to any other goal orientation, so as to verify our theory derived in Sec. IV.
First, based on the single-sphere motion model, the A-F Sphere was grasped and manipulated by the Yale Model O Hand (Fig. 7(a)). We see that with $3$ RRs, the sphere could be rotated to any face. Additionally, the rotations could be done about multiple axes, as opposed to the limited differential motions provided by $2$ RRs as discussed in Sec. IV. The Cheese toy could also be rotated to any desired orientations, as shown in Fig. 7(c), physically justifying the in-hand rotation capability of our RR design, as well as the correctness of our multi-sphere motion model. Furthermore, we found that the form of grasping system used (i.e. Yale Model O Hand and Human hand) made no difference in the ability to fully rotate the grasp objects. Interestingly, as discussed by our motion models and better visualized in the supplementary video, “detours” were often needed for in-hand rotations since the RRs can only provide differential non-holonomic manipulation.
V-B Translational and Rotation Manipulation
@@
After confirming the ability of a full range of reorientation, we moved towards testing the translational manipulation of the RRs. Utilizing the translational portion of the Motion Model (Section IV) we were able to generate translations along all cardinal planes from the base set of actuated RRs. Following the same constraints as the rotational manipulation, so long as the object is stably grasped and does not collide internally with the hand, translations can be generated. Fig. 7(b-d) are examples showing both translations and rotations generated for grasped objects, including a Pringles tube, a Cheese toy, and a cube, to move them within the grasping system. We observed that, similarly to the rotational manipulation, the form of grasping system made no difference whether the RRs were affixed to a human or robot grasping system in terms of manipulation capabilities. So long as the grasping constraints were met, rotations and translations can be generated in sequences to fully manipulate the in-hand object from any pose to any others, as discussed in Sec. IV, although as the same as before, often taking “detours”.
V-C Object Variance
When running the experiments for the RRs, we evaluated how the difference in object geometries affected the manipulation capabilities of the RRs. When affixed to the Yale Model O Hand, we saw that larger objects were capable of being easily manipulated. Objects such as the A-F Sphere and Cube (6) could be smoothly manipulated when held in a power grasp. Sharp corners were difficult to manipulate around as they would push against the contact points when rotated about an axis normal to the set. However, the compliance of the Model O allowed for the RR to complete its manipulation tasks by accounting for the deflection. Objects of smaller geometries (i.e. children’s toys) were more difficult to manipulate as they could only be held in a precision grasp. This loss of stability decreased the ability of the Model O to stably hold the objects during manipulation.
This issue was mirrored when the RR was affixed to a human hand. Objects of smaller geometries (i.e. children’s toys) could be easily manipulated when held in a power grasp, sharp corners increased the difficulty of the grasp, and the compliance of human fingers improved the in-hand manipulation capabilities of the RR. Larger objects, however, could only be grasped in a precision grasp. This increased the difficulty of the in-hand manipulation through a reduction in the stability. Overall, object variance between the relative geometry to the grasping system played a key factor in the ability to successfully and easily manipulate. Objects that were capable of being held in a power grasp were held stably and would not encounter loss of contact or slippage when performing in-hand manipulation.
V-D Limitations of Roller Rings
@@
1) Geometric Constraint: Due to the geometry of the RR and its attachment to the grasping system, objects that are larger than the power grasp task-space or thinner than the formed gap are difficult to manipulate with the base DoF. These objects can be manipulated when held in a precision grasp, but only at a reduced in-hand manipulation capability. Utilizing the palm and fingers as a support in the manipulation of these objects improves the in-hand manipulation, but is not effective across all geometries due to the lack of active surfaces across them.
@@
2) Friction: Because the angle of the CASM was fixed to $30^{\circ}$ to enable a full set of manipulation abilities, the RRs require “slipping and scratching” motions perpendicular to its main motion axes as needed by manipulation actions. Such motions will not only make the system less energy-effective, but also can bring more uncertainties since friction is difficult to model. In addition, the system’s durability can be reduced due to constant “scratching”.
@@
3) Non-Compliance: Due to the RRs not being compliant, contacts between the grasping system and object are difficult to maintain using the RRs alone. Compliant grasping systems such as the Yale Model O and human hands are required in order to perform stable in-hand manipulation. This is because their compliance allows for a higher level of stability when manipulating objects that is not possible with the independent RRs. Compliant mechanism can conform to an object’s geometry in order to adaptively facilitate the manipulation. Without this, affixing the RRs to a non-compliant system will induce difficulties in performing in-hand manipulation as the contacts will be easily broken.
@@ VI Conclusion and Future Work
This paper proposed the design of a wearable Roller Ring, based on the concept of active surfaces, that can be used to enhance the in-hand manipulation capabilities of existing hand designs. Active surface-based motion models were derived to show that active surface-based manipulation in general provides a differential and non-holonomic manipulation model. The motion model also proves that the proposed design is able to provide a complete manipulation skill set including in-hand rotations and translations. Experiments were performed to physically verify the manipulation capabilities of RRs. Different objects of varying sizes and geometries were manipulated to empirically validate the applicability of our design in daily manipulation tasks. In future work, we plan to mechanically make the RRs more compliant, as well as reduce the friction in differential motions. Computationally, based on the derived motion models, we will develop algorithms to provide general and optimized manipulation solutions to more hand designs.

### v4

@@ IV-B Multi-Sphere Motion Model for Arbitrary Object Shapes
Extending from the unit-sphere motion model, we can now model the rotations of arbitrary object shapes as actuated by active surface contacts. We begin with $2$ contacts $p_{a}$ and $p_{b}$. However, this time the contacts are not necessarily on a unit sphere $\mathbb{S}^{2}$. Their distances to the axis of rotation are represented by $r_{a},r_{b}\in\mathbb{R}$. Intuitively, these two contacts are actuating the object as if there were $2$ different contactable spheres, as illustrated in Fig. 5. Thus, we know that the torque generated by these differing unit spheres are not weighted the same given the difference in radii. As such, the torques generated by these two contacts, as shown in Eq. (7), are now scaled by $r_{a},r_{b}$ to account for this difference:
@@
As such, similarly to the derivation of the unit sphere motion model, the torques will balance out when the object is rotating at an equilibrium state. Skipping the intermediate steps, we can obtain the rotation motion model for $N$ active surface contacts on arbitrary objects of non-uniform size:
@@
Based on this multi-sphere motion model, Eq. (8) provides a complete and non-holonomic rotational manipulation solution. However, a major difference now is that the individual angular velocities are additionally weighted by the distances between the contact and the rotation axis, providing another dimension for rotations to be regulated by contact variances.
@@ IV-C Motion Model for In-hand Translations
To achieve a complete in-hand manipulation solution, the final component required in the motion model is the ability to translate the object. The model analysis will focus on pure translations supplementary to the derived rotation models. In practice, however, rotation and translation can happen simultaneously during manipulation in order to perform more efficient actions.
For pure in-hand object translation, the first requirement is that the object’s angular velocity should be zero. Based on our model above, we know that this can be expressed as a motion constraint among all active surfaces: $\omega_{*}=\sum_{i=1}^{N}r_{i}k_{i}\omega_{i}=0$. When this constraint is satisfied, which can be achieved in an infinite number of ways when multiple contacts exist to cancel out each other’s torques, the velocity of the object is purely determined by the linear forces provided by the active surfaces. Let us denote by $v^{*}\in\mathbb{R}^{3}$ the linear velocity of the object, by $v_{i}\in\mathbb{R}^{3}$ and $F_{i}^{t}$ the linear velocity and the tangential force at the $i$-th active surface. The following equation further constrains the object’s translation:
@@
In Eq. (9), the tangential force at each contact is first aligned to the direction of the “slipping or scratching” motion and then projected to the orthogonal direction of the object’s motion by dot product. The zero-sum of all these forces guarantees that the object is purely translating in the direction of $v^{*}$. In practice, although it is not difficult to obtain $\omega_{*}=0$, as discussed above, achieving arbitrary translation motions is challenging due to the additional constraint (Eq. (9)), especially when the number of contacts, $N$, is small. Nonetheless, given our comprehensive solution of in-hand rotation, we will be able to manipulate the object from any pose to any other by sequencing translation and rotations. Thus, a manipulation graph can be constructed to generate such action sequences[34], although the aforementioned “detours” are not avoidable due to the relatively small number of contacts wherein increasing the number of non-colinear contacts can reduce but not entirely remove the degree of detours necessary.
@@ V Experiments
To evaluate the design and motion model of the RR, we constructed a prototype, as seen in Fig. 2 to test its in-hand manipulation capabilities using various objects (Fig. 6). The RRs were affixed to both a Yale Model O hand and a human hand for all manipulation tasks. Each RR had its velocity set to a differing set of values during each experiment to generate differential motions (Sec. IV) and controlled via velocity control teleoperation. No retuning was required between each hand and the same controller was used across all experiments. Execution times were measured by analyzing motion video frames from the start to stop of motions. For each object, the reported time represents the average of six successful trials.
Figure 6: Manipulated experimental objects(Top Row Left to Right: A-F Sphere, Octahedra, Children’s Cheese Toy, Cardboard Tube; Bottom Row Left to Right: A-F Sphere, Cube, Children’s Grape Toy)
@@ V-A Rotational Manipulation
[…]
As discussed by the motion models and visualized in the supplementary video, “detours” were often needed for in-hand rotations since the RRs can only provide differential non-holonomic manipulation. Supplementing more RRs would reduce and/or negate these detours, improving the overall manipulation timing.
Figure 7: Motion Examples of the RR: (a)A-F Sphere reoriented by the RR through all faces when held in a power grasp by the Yale Model O (b) Cardboard tube translated along the Yale Model O when held in a power grasp (c) Cheese toy translated and rotated when held in a precision grasp by the Yale Model O (d) Cube translated and rotated when held in a precision grasp by a human hand
V-B Translational Manipulation
@@
Utilizing the translational portion of the Motion Model (Sec. IV) we were able to generate translations along all cardinal planes from the base set of RRs. Following the same constraints for rotational manipulation, as long as the object did not collide internally and is stably grasped, we found that translations were easily generated. As seen in Fig. 7(b-d), translations were generated for grasped objects of varying geometry such as a Pringles tube, cheese toy, and cube. We observed that, similarly to the rotational manipulation, the type of hand made no difference in terms of manipulation capabilities. Translations can be generated in sequences to fully manipulate the in-hand object from any pose to any others, as discussed in Sec. IV.
[…]
V-C Limitations of Roller Rings
@@
1) Geometric Constraint: Due to the geometry of the RR, objects larger than the hand’s power grasp task-space or thinner than the formed gap are difficult to manipulate. These objects can be manipulated when held in a precision grasp, but with a decrease in stability. Utilizing the palm and fingers as a support improves the in-hand manipulation, but is not effective due to the lack of active surfaces. Future analysis on RR manipulation on certain hand sizes can be performed to better understand its capabilities within this hand-object system.
@@
2) Friction: Because the angle of the CASM was fixed to $30^{\circ}$ to limit the form factor, the RRs require “slipping and scratching” motions orthogonal to its main motion axes for manipulation. Such motions will not only make the system less energy-effective, but can also reduce, the system’s durability due to constant “scratching”. Further experimentation is required to understand and denote the trade-off between lost energy and manipulation timing.
@@
3) Non-Compliance: The RRs are not compliant, thus, contacts between the grasping system and object are difficult to maintain using the RRs alone. Intuitively we can ascertain that because compliant mechanisms can conform to an object’s geometry, the contact generated by the RR will be more stable. Thus, in order to facilitate the manipulation stably, the level of compliance achievable by the RR is directly correlated to the compliance of the underlying gripper. Therefore, RRs on a non-compliant system will encounter difficulties performing in-hand manipulation, as the contacts will be easily broken. By either increasing or having some base level of compliance in the base gripper, the non-compliancy of the RR can be overcome as mirrored in our experimentation of the RR. Compliant grasping systems, such as the Yale Model O or human hand, were required in order to perform stable in-hand manipulation due to their ability to inherently push the RR towards the surface of the object. Because of this, we saw that the human hand had a smaller number of failed trials due to its higher level of compliancy.
@@ VI Conclusion and Future Work
This paper proposed the design of a wearable Roller Ring, utilizing the concept of active surfaces, that can be used to enhance the in-hand manipulation capabilities of existing grasping systems. Active surface-based motion models were derived to show that active surface-based manipulation provides a differential and non-holonomic manipulation model. The motion model also proves that the proposed design is able to provide a complete manipulation skill set, including in-hand rotations and translations. Experiments were performed to physically verify the manipulation capabilities of RRs. Different objects of varying sizes and geometries were manipulated to empirically validate the applicability of our design in daily manipulation tasks. In future work, we plan to mechanically make the RRs more compliant, as well as reduce the differential motion friction. Based on the derived motion models, we will develop algorithms to provide general and optimized manipulation solutions.